OpenStudy (sky7133):

What is the value of expression? 5−8 • 5−4 A. 532 B. 54 C. 5−4 D. 5−12

Nnesha (nnesha):

use `^` for exponents

Nnesha (nnesha):

when we multiply same bases we should `ADD` their exponents \[\huge\rm x^m \times x^n =x^{m+n}\]

OpenStudy (sky7133):

Ok

OpenStudy (sky7133):

So C.

Nnesha (nnesha):

nope \[\huge\rm 5^{-8+(-4)}\] both exponents are negative

Nnesha (nnesha):

so 5^{??} -8+(-4)= ?

OpenStudy (sky7133):

12?

Nnesha (nnesha):

no both numbers are negative so answer would be negative -2-2 isn't 4 -2-2 = `-4`

Nnesha (nnesha):

 1 times -4 = ?